# MoMa: A Modular Deep Learning Framework for Material Property Prediction

### v2

@@ 5 Conclusion & Outlook
In this paper, we present MoMa, a modular deep learning framework for material property prediction. Motivated by the challenges of diversity and disparity, MoMa first trains specialized modules across a wide spectrum of material tasks, constituting MoMa Hub. We then introduce the Adaptive Module Composition algorithm, which facilitates tailored adaptation from MoMa Hub to each downstream task by adaptively composing synergistic modules. Experimental results across 17 datasets demonstrate the superiority of MoMa, with few-shot and continual learning experiments further highlighting its data efficiency and scalability.
@@
Finally, we discuss the prospects of MoMa in driving practical advancements in materials discovery. As an open-source platform enabling materials knowledge modularization and distribution, MoMa offers several key advantages: (1) secure, flexible upload of material modules to MoMa Hub without compromising proprietary data; (2) efficient customization of modules for downstream tasks; (3) enhanced property prediction accuracies, even in low-data scenarios. We envision MoMa facilitating a new paradigm of modular material learning and fostering broader community collaboration toward accelerated materials discovery.
Impact Statement
This paper presents work aimed at advancing the field of materials discovery through innovative machine learning techniques. The potential positive societal impacts include accelerating the discovery of new materials with desirable properties, benefiting industries such as energy, electronics, and manufacturing. However, there are risks associated with the mal-intended use of material knowledge to develop harmful or unsafe materials. To mitigate these risks, it is crucial to ensure that the application of this work adheres to ethical guidelines. While we do not foresee significant negative consequences in the near future, we recognize the importance of responsible usage and oversight in the application of these technologies.
@@ Appendix A Algorithm for Adaptive Module Composition
Algorithm 1 Adaptive Module Composition
1:  Input: MoMa Hub =$\mathcal{H}=\{g_{1},g_{2},\dots,g_{N}\}$, Downstream Task $\mathcal{D}$.
2:  Output: adaptive module $g_{\mathcal{D}}$ for $\mathcal{D}$.
3:  for each module $g_{j}\in\mathcal{H}$ do
4:     Encode the input materials in the training set of $\mathcal{D}$ using $g_{j}$ to obtain $\mathcal{X}^{j}=\{\mathbf{x}_{1}^{j},\mathbf{x}_{2}^{j},\ldots,\mathbf{x}_{m}^{j}\}$.
5:     for each sample $\mathbf{x}_{i}^{j}\in\mathcal{X}^{j}$ do
6:        Compute the predicted label $\hat{y_{i}}^{j}$ for $\mathbf{x}_{i}^{j}$ using $k$NN following Equation 1.
@@
8:  end for
9:  Optimize the module weights $w_{j}$ using cvxpy to minimize the proxy error defined in Equation 2, subject to: $\sum_{j=1}^{N}w_{j}=1$ and $w_{j}\geq 0$. Denote the optimized weights for the $j$-th module as $w_{j}^{*}$.
10:  Compose the final adaptive module $g_{\mathcal{D}}$ by weighted averaging the parameters of the MoMa Hub modules: $g_{\mathcal{D}}=\sum_{j=1}^{N}w_{j}^{*}g_{j}$
11:  Return: The composed module $g_{\mathcal{D}}$.
Appendix B Experimental Details
@@
In this section, we provide more experimental details of MoMa regarding the datasets, implementation, baselines, and the continual learning setting.
B.1 Dataset Details
@@
We primarily adopt the dataset setup proposed by Chang et al. (2022). Specifically, we select 35 datasets from Matminer (Ward et al., 2018) for our study, categorizing them into 18 high-resource material tasks, with sample sizes ranging from 10,000 to 132,000 (an average of 35,000 samples), and 17 low-data tasks, with sample sizes ranging from 522 to 8,043 (an average of 2,111 samples).
The high-resource tasks are utilized for training the MoMa Hub modules, as their larger data volumes are likely to encompass a wealth of transferrable material knowledge. A detailed introduction of these MoMa Hub datasets is included in Table 2.
The low-data tasks serve as downstream datasets to evaluate the effectiveness of MoMa and its baselines. This setup mimics real-world materials discovery scenarios, where downstream data are often scarce. To ensure robust and reliable comparison results, we exclude two downstream datasets with exceptionally small data sizes (fewer than 20 testing samples) from our experiments, as their limited data could lead to unreliable conclusions. A detailed introduction is included in Table 3.
Following Chang et al. (2022), all datasets are split into training, validation, and test sets with a ratio of 7:1.5:1.5. For the downstream low-data tasks, the splitting is performed randomly for 5 times to ensure the stability of evaluation.
B.2 Implementation Details of MoMa
Network Architecture
We now introduce the network architecture of MoMa modules. The JMP (Shoghi et al., 2023) backbone is directly taken as the full module parametrization. JMP is pre-trained on $\sim$ 120 million DFT-generated force-field data across large-scale datasets on catalyst and small molecules. JMP is a 6-layer GNN model with around 160M parameters which is based on the GemNet-OC architecture (Gasteiger et al., 2022). Note that MoMa is backbone-agnostic. JMP is selected due to its comprehensive strength across a wide range of molecular and crystal tasks, which allows us to seamlessly conduct the continual learning experiments. We leave the extrapolation of MoMa to other architectures as future work.
For the adapter module, we follow the standard implementation of adapter layers (Houlsby et al., 2019). Specifically, we insert adapter layers between each layer of the JMP backbone. Each layer consists of a downward projection to a bottleneck dimension and an upward projection back to the original dimension.
@@ B.2 Implementation Details of MoMa
For the training of JMP backbone, we mainly follow the hyper-parameter configurations in Shoghi et al. (2023), with slight modifications to the learning rate and batch size. During the module training stage of MoMa, we use a batch size of 64 and a learning rate of 5e-4 for 80 epochs. During downstream fine-tuning, we adopt a batch size of 32 and a learning rate of 8e-5. We set the training epoch as 60, with an early stopping patience of 10 epochs to prevent over-fitting. We adopt mean pooling of embedding for all properties since it performs significantly better than sum pooling in certain tasks (e.g. band gap prediction), which is consistent with findings in Shoghi et al. (2023).
For the adapter modules, we employ BERT-style initialization (Devlin, 2018), with the bottleneck dimension set to half of the input embedding dimension.
For the Adaptive Module Composition (AMC) algorithm, we set the number of nearest neighbors ($K$ in Equation 1) to 5. For the optimization problem formulated in Equation 3, we utilize the CPLEX optimizer from the cvxpy package (Diamond et al., 2014). AMC is applied separately for each random split of the downstream tasks to avoid data leakage.
@@
Experiments are conducted on NVIDIA A100 80 GB GPUs. During the module training stage, training time ranges from 30 to 300 GPU hours, depending on the dataset size. While this training process is computationally expensive, it is a one-time investment, as the trained models are stored in MoMa Hub as reusable material knowledge modules. Downstream fine-tuning requires significantly less compute, ranging from 2 to 8 GPU hours based on the dataset scale. The full module and adapter module require similar training time; however, the adapter module greatly reduces memory consumption during training.
B.3 Baseline Details
The CGCNN baseline refers to fine-tuning the CGCNN model (Xie & Grossman, 2018) separately on 17 downstream tasks. Conversely, MoE-(18) involves training individual CGCNN models for each dataset in MoMa Hub and subsequently integrating these models using mixture-of-experts (Jacobs et al., 1991; Shazeer et al., 2016). For the baseline results of CGCNN and MoE-(18), we adopt the open-source codebase provided by Chang et al. (2022) and follow the exactly same parameters as reported in their papers for the result duplication.

### v3

@@ 5 Conclusion
In this paper, we present MoMa, a simple modular learning framework for material property prediction. Motivated by the challenges of diversity and disparity in materials, MoMa first trains specialized modules across a wide spectrum of material tasks, constituting MoMa Hub. We then introduce the Adaptive Module Composition algorithm, which facilitates tailored adaptation from MoMa Hub to each downstream task by adaptively composing synergistic modules. Experimental results across 17 datasets demonstrate the superiority of MoMa, with few-shot and hub-scaling experiments further highlighting its data efficiency and scalability.
Limitations and Future Work. The current scope of our study is limited to crystalline and organic materials. Future work includes expanding MoMa Hub with modules for a wider range of material data and prediction tasks, and examining how MoMa scales with hundreds or thousands of modules, which may yield deeper insights into the modularity of materials knowledge.
@@
Broader Impact. As an open-source platform for modularizing and distributing materials knowledge, MoMa enables secure sharing of modules without exposing proprietary data, efficient customization for downstream tasks, and improved prediction accuracy even in low-data scenarios. We envision MoMa fostering a new paradigm of modular material learning and driving broader community collaboration toward accelerated materials discovery.
Acknowledgments
The authors would thank Junwei Yang, Ziyao Cao, Fanyou Meng, and Yuxuan Song for their valuable feedback on the paper. We also thank the anonymous reviewers for reviewing the draft. This work is supported by the National Science and Technology Major Project (2022ZD0117502), the Natural Science Foundation of China (Grant No. 62376133, 62406170), and Wuxi Research Institute of Applied Technologies, Tsinghua University under Grant 20242001120.
@@ Appendix A Algorithm for Adaptive Module Composition
Algorithm 1 Adaptive Module Composition (AMC)
1:Input: MoMa Hub $\mathcal{H}=\{g_{j}\}_{j=1}^{N}$, Downstream training set $\mathcal{D}=\{(X_{i},y_{i})\}_{i=1}^{m}$.
2:Output: Composed module $g_{\mathcal{D}}$.
3: {1. Module Prediction Estimation}
4:for $j=1\to N$ do
5:  Generate embeddings $\mathcal{X}^{j}\leftarrow\{g_{j}(X_{i})\mid(X_{i},y_{i})\in\mathcal{D}\}$.
6:  Compute prediction vector $\hat{\mathbf{Y}}^{j}=(\hat{y}_{1}^{j},\dots,\hat{y}_{m}^{j})$ via leave-one-out $k$-Nearest Neighbors.
@@
8: {2. Module Weight Optimization}
9: Let $\mathbf{Y}=(y_{1},\dots,y_{m})$ be the vector of true labels from $\mathcal{D}$.
10: Find optimal weights $\bm{w}^{*}=(w_{1}^{*},\dots,w_{N}^{*})$ by solving the convex optimization problem:
[…]
12:  subject to: $\sum_{j=1}^{N}w_{j}=1$ and $w_{j}\geq 0$ for all $j\in\{1,\dots,N\}$.
13: {3. Module Composition}
[…]
15:Return $g_{\mathcal{D}}$
Appendix B Theoretical Justification and Error Analysis for AMC
In this section, we provide a formal analysis of the kkNN-based proxy error used in AMC. Specifically, we show that the kkNN proxy risk $R_{proxy}(w)$ serves as an upper bound for the fine-tuning risk $R_{FT}(w)$ (subject to approximation errors). Consequently, minimizing the empirical approximation of $R_{proxy}(w)$ tightens this bound, thereby providing theoretical justification for using the proxy risk to control the risk of the subsequently fine-tuned model.
B.1 Definitions
Let $\theta_{i}$ denote the parameters of the $i$-th module and define its representation of input $x$ as $g_{i}(x):=g(\theta_{i};x)$. Given weights $w=(w_{1},\dots,w_{N})\in\Delta_{N-1}$, define the merged module in parameter space and its representation by
$\theta_{w}:=\sum_{i=1}^{N}w_{i}\,\theta_{i},\qquad g_{w}(x):=g(\theta_{w};x).$
Let the Bayes regressors associated with each representation be
$m_{i}(x):=\mathbb{E}[\,Y\mid g_{i}(X)=g_{i}(x)\,],\quad m_{w}(x):=\mathbb{E}[\,Y\mid g_{w}(X)=g_{w}(x)\,].$
We define the following risk terms:
  • Representation Bayes Risk: $R^{*}(w):=\mathbb{E}\big[(m_{w}(X)-Y)^{2}\big]$
  • Fine-tuning Risk: $R_{\mathrm{FT}}(w):=\mathbb{E}\big[(\hat{y}_{\mathrm{FT}}(X;w)-Y)^{2}\big]$
  • Proxy Risk using kNN: $R_{\mathrm{proxy}}(w):=\mathbb{E}\big[(\hat{y}_{\mathrm{proxy}}(X;w)-Y)^{2}\big]$
  • Bayes Ensemble Risk: $R_{\mathrm{ens}}(w):=\mathbb{E}\big[(m_{\mathrm{ens}}(X;w)-Y)^{2}\big]$, where $m_{\mathrm{ens}}(x;w):=\sum_{i=1}^{N}w_{i}\,m_{i}(x)$.
Remark (distribution vs empirical proxy objective).
Practically AMC optimizes the empirical proxy error
$E_{\mathcal{D}}(w)=\frac{1}{m}\sum_{k=1}^{m}\Big(\sum_{i=1}^{N}w_{i}\,\hat{y}_{i,k}-y_{k}\Big)^{2},$
while analysis here is stated for the distribution-level proxy risk $R_{\mathrm{proxy}}(w)$. Under standard generalization results for squared-loss regression and mild capacity control on the family $\{\hat{y}_{\mathrm{proxy}}(\cdot;w):w\in\Delta_{N-1}\}$, the empirical proxy error $E_{\mathcal{D}}(w)$ concentrates around $R_{\mathrm{proxy}}(w)$; hence we treat $E_{\mathcal{D}}(w)$ in practice as a finite-sample approximation of $R_{\mathrm{proxy}}(w)$.
B.2 Preliminaries and Assumptions
We recall a standard non-parametric regression result on the consistency of kNN estimators.
Lemma 1 (Universal kNN $L_{2}$-Consistency (Stone, 1977; Devroye et al., 1994)).
For each upstream task $i$, consider a training dataset of size $n$, and let $\hat{y}_{n}^{(i)}$ be the k-nearest neighbor regressor in the feature space $g_{i}(x)$, where the number of neighbours $k=k_{n}$ satisfies $k_{n}\to\infty,\quad k_{n}/n\to 0$. Define the $L_{2}$-estimation error
$\varepsilon^{(i)}_{\mathrm{kNN}}(n):=\mathbb{E}\big[(\hat{y}_{n}^{(i)}(X)-m_{i}(X))^{2}\big].$
Then the kNN regression estimator satisfies the exact risk decomposition
$\mathbb{E}\big[(\hat{y}_{n}^{(i)}(X)-Y)^{2}\big]=\mathbb{E}\big[(m_{i}(X)-Y)^{2}\big]+\varepsilon^{(i)}_{\mathrm{kNN}}(n),$
and moreover $\varepsilon^{(i)}_{\mathrm{kNN}}(n)\;\to\;0$ as $n\to\infty.$
Remark on Lemma 1.
Our implementation of AMC uses a finite-$K$ kNN ensemble with cosine similarity in the learned representation space. This differs from the asymptotic setting of Lemma 1 (which requires $k_{n}\to\infty$ and $k_{n}/n\to 0$), but it serves as a computationally efficient finite‐sample approximation. Lemma 1 is used as a conceptual tool to justify the kNN-based proxy objective; we do not claim a full universal consistency result for our exact finite-sample variant.
Assumption 1 (Fine-tuning stability).
There exists $\varepsilon_{\mathrm{opt}}\geq 0$ such that for all $w\in\Delta_{N-1}$,
$R_{\mathrm{FT}}(w)\;\leq\;R^{*}(w)+\varepsilon_{\mathrm{opt}}.$
That is, after fine-tuning on top of $g_{w}$, the resulting predictor is within $\varepsilon_{\mathrm{opt}}$ of the optimal predictor defined on the same representation (McNamara and Balcan, 2017; Mehra and others, 2024).
Assumption 2 (Representation closeness).
There exists $\delta>0$ and a high-probability subset $\mathcal{X}_{0}\subseteq\mathcal{X}$ such that for all modules $i$ and all $x\in\mathcal{X}_{0}$,
$\|g_{i}(x)-g_{w}(x)\|\;\leq\;\delta.$
This models the empirical observation that independently fine-tuned modules from the same pre-trained model can often be aligned into a shared, approximately convex basin in parameter space, leading to similar internal representations. Recent work on mode connectivity (Frankle et al., 2020; Entezari et al., ) and cross-task linearity (Zhou et al., 2024) supports this assumption, which we adopt here as a structural modeling assumption rather than a general theorem.
Assumption 3 (stability of latent regressor).
Let $g$ and $g^{\prime}$ denote any two modules among $\{g_{i}:i\in[N]\}$ and the merged module $g_{w}$. Let their associated Bayes regressors be
$m(x):=\mathbb{E}[\,Y\mid g(X)=g(x)\,],\qquad m^{\prime}(x):=\mathbb{E}[\,Y\mid g^{\prime}(X)=g^{\prime}(x)\,].$
There exists $L>0$ and a high-probability subset $\mathcal{X}_{0}\subseteq\mathcal{X}$ such that if
$\|g(x)-g^{\prime}(x)\|\leq\delta\qquad\forall x\in\mathcal{X}_{0},$
then the regressors are close in expectation:
$\mathbb{E}\big[\,|m(X)-m^{\prime}(X)|\,\mathbf{1}\{X\in\mathcal{X}_{0}\}\big]\;\leq\;L\,\delta.$
We also assume $|Y|\leq B$ almost surely.
Remark on Assumption 3.
Assumption 3 formalizes a semantic smoothness condition in the learned representation: inputs that are mapped to nearby latent points (i.e., with close $g(x)$ and $g^{\prime}(x)$) are required to have similar predictive behavior for $Y$ on average. This can be viewed as a Lipschitz-type regularity assumption on the Bayes regressors in representation space, and is consistent with the common inductive bias in deep learning that well-trained encoders should support target functions which do not change abruptly under small perturbations of the latent features.
B.3 Risk Transfer Analysis
Step 1: kNN Ensemble Approximates the Bayes Ensemble.
We compare the proxy risk to the Bayes ensemble risk:
$\big|R_{\mathrm{proxy}}(w)-R_{\mathrm{ens}}(w)\big|$ $=\big|\mathbb{E}\big[(\hat{y}_{\mathrm{proxy}}-Y)^{2}-(m_{\mathrm{ens}}-Y)^{2}\big]\big|$
$=\big|\mathbb{E}\big[(\hat{y}_{\mathrm{proxy}}-m_{\mathrm{ens}})(\hat{y}_{\mathrm{proxy}}+m_{\mathrm{ens}}-2Y)\big]\big|$
$\leq 4B\,\mathbb{E}\big[\,|\hat{y}_{\mathrm{proxy}}-m_{\mathrm{ens}}|\big]\qquad\text{(by bounded label in Assumption 3)}$
Next,
$\mathbb{E}\big[\,|\hat{y}_{\mathrm{proxy}}-m_{\mathrm{ens}}|\big]$ $=\mathbb{E}\Big[\,\Big|\sum_{i=1}^{N}w_{i}\hat{y}^{(i)}-\sum_{i=1}^{N}w_{i}m_{i}\Big|\Big]$
$\;\leq\;\sum_{i=1}^{N}w_{i}\,\mathbb{E}\big[\,|\hat{y}^{(i)}-m_{i}|\big]$
$\;\leq\;\sum_{i=1}^{N}w_{i}\sqrt{\mathbb{E}\big[\,(\hat{y}^{(i)}-m_{i})^{2}\big]}$
$=\sum_{i=1}^{N}w_{i}\sqrt{\varepsilon^{(i)}_{\mathrm{kNN}}(n)}\qquad\text{(by Assumption 1)}$
$\;\leq\;\sqrt{\sum_{i=1}^{N}w_{i}\,\varepsilon^{(i)}_{\mathrm{kNN}}(n)}\qquad\text{(by Jensen's inequality)}$
Define the weighted kNN error
$\bar{\varepsilon}_{\mathrm{kNN}}(n;w):=\sum_{i=1}^{N}w_{i}\,\varepsilon^{(i)}_{\mathrm{kNN}}(n),\qquad\epsilon_{1}(n;w):=4B\sqrt{\,\bar{\varepsilon}_{\mathrm{kNN}}(n;w)\,}.$
Then we obtain
$\big|R_{\mathrm{proxy}}(w)-R_{\mathrm{ens}}(w)\big|\;\leq\;\epsilon_{1}(n;w).$
Step 2: Bayes Ensemble Approximates the Merged Bayes Predictor.
We compare $m_{\mathrm{ens}}(x;w)$ and $m_{w}(x)$ using the triangle inequality,
$\mathbb{E}\!\left[\,\big|m_{\mathrm{ens}}(X;w)-m_{w}(X)\big|\,\mathbf{1}\{X\in\mathcal{X}_{0}\}\right]$ $=\mathbb{E}\!\left[\,\Big|\sum_{i=1}^{N}w_{i}m_{i}(X)-m_{w}(X)\Big|\,\mathbf{1}\{X\in\mathcal{X}_{0}\}\right]$
$\leq\sum_{i=1}^{N}w_{i}\,\mathbb{E}\!\left[\,|m_{i}(X)-m_{w}(X)|\,\mathbf{1}\{X\in\mathcal{X}_{0}\}\right]$
$\leq\;L\,\delta\qquad\text{(by Assumption2 and 3)}$
Then
$\big|R_{\mathrm{ens}}(w)-R^{*}(w)\big|$ $=\Big|\mathbb{E}\big[(m_{\mathrm{ens}}(X;w)-Y)^{2}-(m_{w}(X)-Y)^{2}\big]\Big|$
$=\Big|\mathbb{E}\big[(m_{\mathrm{ens}}(X;w)-m_{w}(X))(m_{\mathrm{ens}}(X;w)+m_{w}(X)-2Y)\big]\Big|$
$\leq 4B\,\mathbb{E}\!\left[\,\big|m_{\mathrm{ens}}(X;w)-m_{w}(X)\big|\right]$
$\leq 4B\,L\,\delta+\text{(small error term on }\mathcal{X}_{0}^{c}).$
Absorbing the small-probability contribution from $\mathcal{X}_{0}^{c}$ into the constant, we obtain
$\big|R_{\mathrm{ens}}(w)-R^{*}(w)\big|\;\leq\;C\,\delta,$
where $C:=4BL$.
B.4 Main Transfer Bound and Guarantee
Proposition 1.
Under Assumptions 1–3 and Lemma 1, for any $w\in\Delta_{N-1}$,
$R_{\mathrm{FT}}(w)\;\leq\;R_{\mathrm{proxy}}(w)+C\,\delta+\varepsilon_{\mathrm{opt}}+\epsilon_{1}(n;w),$
where $C=4BL$ and $\epsilon_{1}(n;w)=4B\sqrt{\,\bar{\varepsilon}_{\mathrm{kNN}}(n;w)\,}$.
Proof. From Step 1 we have
$R_{\mathrm{proxy}}(w)\;\geq\;R_{\mathrm{ens}}(w)-\epsilon_{1}(n;w),$
and from Step 2
$R_{\mathrm{ens}}(w)\;\geq\;R^{*}(w)-C\,\delta.$
Hence
$R^{*}(w)\;\leq\;R_{\mathrm{proxy}}(w)+C\,\delta+\epsilon_{1}(n;w).$
Combining with Assumption 1 yields
$R_{\mathrm{FT}}(w)\;\leq\;R^{*}(w)+\varepsilon_{\mathrm{opt}}\;\leq\;R_{\mathrm{proxy}}(w)+C\,\delta+\varepsilon_{\mathrm{opt}}+\epsilon_{1}(n;w).$
Combining this result with the optimality of $\hat{w}$ for the proxy objective yields the following theorem.
Theorem 1.
Let $\hat{w}\in\arg\min_{w\in\Delta_{N-1}}R_{\mathrm{proxy}}(w)$ be a minimizer of the proxy risk. Under Assumptions 1–3 and Lemma 1, the fine-tuning risk of the merged encoder with weights $\hat{w}$ satisfies
[…]
where C=4BLC=4B\,L and $\epsilon_{1}(n;\hat{w})=4B\sqrt{\,\sum_{i=1}^{N}\hat{w}_{i}\,\varepsilon^{(i)}_{\mathrm{kNN}}(n)\,}$.
Appendix C Experimental Details
@@
Here we provide more experimental details regarding the datasets, baselines, and implementation.
C.1 Dataset Details
@@
We primarily adopt the dataset setup proposed by Chang et al. (2022). Specifically, we select 35 datasets from Matminer (Ward et al., 2018) for our study, categorizing them into 18 high-resource material datasets, with sample sizes ranging from 10,000 to 132,000 (an average of 35,000 samples), and 17 low-data datasets, with sample sizes ranging from 522 to 8,043 (an average of 2,111 samples).
The high-resource datasets are utilized for training the MoMa Hub modules, as their larger data volumes are likely to encompass a wealth of transferrable material knowledge. A detailed introduction of these MoMa Hub datasets is included in Table 4.
The low-data datasets serve as downstream tasks to evaluate the effectiveness of MoMa and its baselines. A detailed introduction is included in Table 5. This setup mimics real-world materials discovery scenarios, where downstream data are often scarce. Compared to the benchmark in Chang et al. (2022), we exclude two low-data datasets with exceptionally small data sizes (fewer than 20 test samples) from our experiments, as their limited data could lead to unreliable conclusions.
Following Chang et al. (2022), all datasets are split into training, validation, and test sets with a ratio of 7:1.5:1.5. For the downstream low-data datasets, we follow the exact splitting provided by Chang et al. (2022) to ensure a fair comparison.
C.2 Implementation Details of MoMa
Module Architecture Details
We now introduce the architectural details of MoMa modules. Across all our experiments in the main text, the JMP (Shoghi et al., 2024) backbone is adopted due to its comprehensive strength across a wide range of molecular and crystal tasks. JMP is pre-trained on $\sim$ 120 million DFT-generated force-field data across large-scale datasets on catalyst and small molecules. It is a 6-layer GNN model with around 160M parameters which is based on the GemNet-OC architecture (Gasteiger et al., 2022). Note that MoMa is backbone-agnostic and we include results with the Orb model (Neumann et al., 2024) in Section 4.3.
For the full module parametrization, we exclude the output layer and treat the entire GNN backbone as a single module. For the adapter components, we follow the standard implementation of adapter layers (Houlsby et al., 2019). Specifically, an adapter layer is inserted between every two layers of the JMP backbone. Each adapter consists of a downward projection to a bottleneck dimension, followed by an upward projection back to the original dimension. We adopt BERT-style initialization (Devlin, 2018), with the bottleneck dimension set to half of the input embedding dimension. Note that the merging process for adapters is performed in a layer-wise manner. For each backbone layer containing adapters, we compute a weighted average of the parameters from all selected adapter modules. A single scalar weight for each module, determined by AMC, is applied uniformly across all adapter layers belonging to that module.
@@ C.2 Implementation Details of MoMa
For the training of JMP backbone, we mainly follow the hyper-parameter configurations in Shoghi et al. (2024), with slight modifications to the learning rate and batch size. During the module training stage of MoMa, we use a batch size of 64 and a learning rate of 5e-4 for 80 epochs. During downstream fine-tuning, we adopt a batch size of 32 and a learning rate of 8e-5. We set the training epoch as 60, with an early stopping patience of 10 epochs to prevent over-fitting. We adopt mean pooling of embedding for all properties since it performs significantly better than sum pooling in certain tasks (e.g. band gap prediction), which echos the findings in Shoghi et al. (2024).
For the Adaptive Module Composition (AMC) algorithm, we set the number of nearest neighbors ($K$ in Eq. 1) to 5. For the optimization problem formulated in Eq. 3, we utilize the CPLEX optimizer from the cvxpy package (Diamond et al., 2014). AMC is applied separately for each random split of the downstream tasks to avoid data leakage.
@@
Experiments are conducted on NVIDIA A100 80 GB GPUs. During the module training stage, training time ranges from 30 to 300 GPU hours, depending on the dataset size. While this training process is computationally expensive, it is a one-time investment, as the trained models are stored in MoMa Hub as reusable material knowledge modules. Downstream fine-tuning requires significantly less compute, ranging from 2 to 8 GPU hours based on the dataset scale. The full module and adapter module require similar training time; however, the adapter module greatly reduces memory consumption during training. The time cost of AMC is discussed in Section D.3.
C.3 Baseline Details
The CGCNN baseline refers to fine-tuning the CGCNN model (Xie and Grossman, 2018) separately on 17 downstream tasks. Conversely, MoE-(18) involves training individual CGCNN models for each dataset in MoMa Hub and subsequently integrating these models using mixture-of-experts (Jacobs et al., 1991; Shazeer et al., 2016). For the baseline results of CGCNN and MoE-(18), we reproduce the results with the open-source codebase provided by Chang et al. (2022) and follow the exactly same hyper-parameters as reported in their papers.
For UMA, we fine-tune the UMA-Medium checkpoint (the largest open-sourced UMA model) on each dataset. To determine the batch size, we follow the max-atom batching strategy from the original UMA paper and set the maximum atoms per batch to 200, which ensures consistent memory usage across systems of varying sizes. All remaining hyperparameters (e.g., learning rate, number of epochs) follow the configurations used in JMP baselines.